# InstructPipe: Building Visual Programming Pipelines with Human Instructions

### v1

@@ 6.1. Study Design
In the user evaluation, we aimed to investigate how the interface condition (with InstructPipe and without InstructPipe; the independent variable) affects the user experience and behaviors (the dependent variable). We will refer to these two interface conditions as “InstructPipe” and “Visual Blocks” in the following content. Figure 10 visualizes the complete study flow. In each condition, participants completed the two pipelines with counterbalance (referred to as Task 1 and Task 2 in Figure 10).
We carefully designed this study to ensure a fair study that can be completed with reasonable efforts. The following elaborates how we made two important decisions that affected the rigorousness of our study:
  • Two controlled pipelines. Our user evaluation focuses on two controlled pipelines. While we acknowledge more pipelines (e.g., four, six, or even more, instead of two in our design) can enhance the solidness of the study, such design also dramatically increases the required groups of the user study session. For example, counterbalancing four controlled pipelines with two interface conditions results in $4!\times 2=48$ study orders. If there are four participants for each unique order, the whole study requires $48\times 4=192$ participants, which is far more than a standard number in HCI system papers (i.e., recruiting 10 - 20 participants for evaluation). We believe two pipelines fit the best in our case since they require $2!\times 2\times 4=16$ participants.
  • Pipeline selection. With the limited number of pipelines we can choose in the study, it becomes highly important how we select the two pipelines. There are two critical factors we considered in the pipeline selection process: representativeness and diversity. The representativeness implies that the selected pipelines should represent the averaged performance of InstructPipe. The diversity further suggests that the selected pipelines should provide various experiences to simulate the actual use scenarios in which InstructPipe may work well in some pipelines but not always. By following this guideline, we first selected four candidates, and the final decision was made after a pilot study with one participant to test the level of pipeline difficulty. The two resultant pipelines are composed of eight nodes with seven edges and six nodes with six edges, respectively. Using the instructions from two authors, the averaged ratio of human interactions in these two pipelines are $27.8\%$ and $5.2\%$, respectively. See Appendix C for more details of the pipelines.
@@ 6.2. Procedure
The study starts with a 10-15 minute hands-on training of both conditions. The training included 1) all the Visual Blocks interactions needed to complete the subsequent steps of the experiment and 2) all the nodes that participants need to use for creating the pipelines they are assigned in the main sessions. Participants were also encouraged to try building a pipeline independently and to ask questions.
After the training, participants progressed to an unmoderated session where they were asked to build pipelines under the given conditions. We verbally described the pipelines to participants as below, and participants are not allowed to read our scripts:
@@
During the task, participants were allowed to consult with us for technical help. If participants were unable to make progress, we provided hints. We provided such support to ensure that each participant spent a reasonable amount of time on each task and had sufficient time for the following tasks. As shown in the results (section 6.5.2), we provided far more support in the “Visual Blocks” condition. Therefore, we argued that this is a fair moderation in the study because it favors our comparative conditions (i.e., it makes it easier for users to finish the pipeline under the “Visual Blocks” condition). As an optional extension to the study, eight participants were offered an open-ended pipeline creation, where participants prototype their own ideas with InstructPipe. This optional section was offered based on the progress of the participant in the previous sections and time constraints so that the study duration was controlled within the time we guaranteed in our recruitment process.
After trying all pipeline-condition combinations, participants answered open-ended questions in a semi-structured interview. The interview script is available in the appendix §D. Participants provided their general outlook of each condition, listed pros and cons, identified potential future use cases, and critiqued the user interface for future improvements.
In total, participants spent 55-65 minutes in the study.
@@ 6.3. Participants
We recruited 16 participants from an internal participant pool at Google. A diverse variety of job profiles were represented, but no software engineers were included. See Table (Table 2) for a full breakdown. Note that we intentionally recruited novice users, as we envision them as intended users of InstructPipe. The criteria for selecting the novice users was based on self-evaluation ratings of the below prompts:
  • Please provide a self-evaluation of your programming experience
  • Please provide a self-evaluation of your machine learning (ML) skillset
@@ 6.4.1. Task Completion Time
Back-end logs were used to collect timestamps for starting and ending events. Then, completion times for each condition were calculated per task for each participant.
@@ 6.5. Results
Figure 11. Raw-TLX results. The statistic significance is annotated with $*{}^{*}$, ${}^{**}$, or ${}^{***}$ (representing $p$¡.05, $p$¡.01, and $p$¡.001, respectively).
Table 3. Task completion time and the number of human interactions in the user study (N=16).
@@ 6.5.1. InstructPipe Reduces Users’ Workload
Figure 11 further visualizes the results of users’ perceived workload in six sub-scales. The Wilcoxon signed ranks test revealed significant differences on five sub-scales, all but “Mental Demand” (see section 7.2 for more explanations and discussion). Furthermore, the test indicates that all participants unanimously vote InstructPipe provides lower or equal workload on the subscales of “Physical Demand”, “Temporal Demand”, “Performance” and “Effort” ($W=0$). These quantitative results, with both objective and subjective metrics, demonstrate that InstructPipe can effectively reduce users’ workload during the pipeline creation process.
@@ 6.5.2. On-boarding Support of Visual Programming
P1, P5, and P9 explicitly mentioned that there is a “learning curve” in the visual programming system for beginners, which validates our statements in section 1 that motivates this project.
“There is a learning curve to it (using the visual programming system) for sure, because you have to, like, read each node carefully” [P1]
P1’s comment matches with our observation of participants’ behaviors during the study. In the Visual Blocks condition, we observed that people were more easily stuck in their creative purposes, which required our support [^4] In the whole study, we encouraged the users to talk to us while keeping their focus on their task. While we offer explicit hints only when we consider a participant is stuck in an issue, it is still difficult to strictly define what conversation is “ providing hints” and count them quantitatively. Therefore, we cannot fairly report quantitative data for “the number of hints” in both conditions. Here, we faithfully report this finding based on our observation and experience. . Typical supports include 1) guiding participants if we notice they go too far away from the correct pipeline and 2) reminding them of an important node for the pipeline, although we introduced all the necessary nodes in our training session.
@@ 6.5.3. Integration into The Existing Workflow
InstructPipe is a feature available in Visual Blocks. In the interviews, participants particularly expressed their strong appreciation of this design as an AI assistant that enhances, instead of completely replacing, the existing user workflow (i.e., the creative process purely using the node-graph editor):
@@
While most participants shared similar opinions with P6, P15 raised one concern about such integration. The concern lies in the mental experience of the state that transits from the prompting task to the node-graph editing process:
“Rather than fixing it (the generated pipeline) on my own [on the node-graph editor], I would have gone back and changed my prompt … I’m pretty sure I could have gotten it closer … because I just spent so much time figuring out what the prompt should be. That’s kind of like already where my brain was and I knew that something was wrong there (the prompt), but I would have to switch over to the other mode (visual programming) of figuring out what was wrong on the pipeline.” [P15]
@@ 6.5.4. Use Scenarios: Accessible ML Prototyping and Education
In the open-ended session, we observed participants could efficiently utilize InstructPipe to prototype a pipeline for various daily life or business purposes. For example, P14 tried InstructPipe with “summarize real estate price increase in San Diego California over 2023”. Compared to using LLM chatbots, InstructPipe helps the user quickly build a more explainable pipeline in which the user can track (or modify) the information resources. P4 prototyped an interactive VQA app by “Describe the product in the camera”. P13 further shared his thoughts on how this rapid and accessible prototyping experience can support future business:
“It (InstructPipe)’s going to facilitate the prototype building for PMs (Product Managers) … I was a PM … Back then … My biggest fear is to code … I have lots of ideas, but my challenge is how to translate an idea into the technical world and see a prototype. I think that this app expedites me in that process a lot.” [P13]
Another emerging theme was regarding educating kids on programming as explained in the following:
@@ 6.5.5. Limitations and Future Directions
Across the study sessions, we consistently observed a specific user behavior pattern: participants typically paused their pace when a generated pipeline appeared in the workspace. At these times, some participants used soliloquy, as in saying “Let me see”, while others kept a focused stare on the workspace. These human behaviors strongly indicate that the participants were engaging in deep, contemplative thought.
This observation suggests that participants typically need to perceive a pipeline that appears in the workspace. Such a sense-making process brings new challenges to their creative process:
@@
Additionally, we observed that several participants spent more time crafting their prompts than the others. P15 is the participant who spent the most time writing the prompt. The following comments explain the reason, and how the prompting process causes extra mental workload to him/her/them.
@@ 7.1. Human-AI Collaboration in Prototyping Open-ended ML Pipelines
Our technical evaluation (section 5.4) shows that InstructPipe reduces the number of user interactions to 18.9 % (±$\pm 20.3\%$). There are two implications from the results.
@@
Such takeaways differ from existing findings that show LLMs can achieve full automation of ML inference (Surís et al., 2023; Gupta and Kembhavi, 2023). The main reason is that existing work built their ad hoc solutions for target use scenarios, respectively. In contrast, InstructPipe covers a wide range of mainstream ML models (section 3.2) and aims for an open-ended use case. Our results show that LLMs (i.e., GPT-3.5-turbo) still fail to write robust code with prompt engineering techniques. As we mentioned in section 4.4, the main attribute of such failure is that LLMs tend to invent nodes (i.e., the node type in the pseudocode) that do not exist in our node library, which causes execution error. Note that we explicitly instructed the LLM that “introducing new nodes is not permitted” in the system prompt for the Code Writer module (Figure 7). We also observed similar issues when we combined all the official prompts of four scenarios in VisProg (Gupta and Kembhavi, 2023).
While the execution error in existing work indicates the failure of the full solution, InstructPipe renders the executable components of the generation pipeline in a node-graph editor. This facilitates a collaboration protocol between humans and AI in which humans do not need to start a project from scratch. Results in section 6.5.3 revealed that a wide range of participants appreciated this design of human-AI collaboration. We believe the main reason behind this observation is that most non-expert users prefer working on a task from a template (although sometimes it is not perfect) rather than starting the whole project from scratch. The visual programming system provides non-experts with a novice-friendly interface to work with AI in a visual platform. In this manner, “users can do what matters most – building software by letting AI do the redundant work” (Zhang et al., 2023a).
7.2. Human-AI Collaboration in Prototyping Open-ended ML Pipelines
@@
While most participants are positive about our integration of InstructPipe into the existing user workflow of Visual Blocks, results from the user evaluation also reveal several challenges in such human-AI collaboration. Results in section 6.5.1 show that InstructPipe failed to significantly reduce novice users’ mental demand. We summarized its major causes into three aspects.
@@
  • Instruction. P15’s comment in section 6.5.5 summarizes the first aspect that causes mental burden. Although the “instruction-to-pipeline” process is fast and seems effortless, the process of framing a prompt is one factor that may overwhelm users, especially those who are more accustomed to visual thinking. In essence, an instructor of an LLM system (the user of InstructPipe) should be clear about the problem they want to solve and preferably what pipeline they want, which causes a mental burden to the user. Zamfirescu-Pereira et al.’s study shows that non-experts may not prompt LLMs well (Zamfirescu-Pereira et al., 2023), which further explains why non-expert participants might find the instruction process mentally demanding.
@@
  • Perception. The integration of LLM supports into the visual programming interface enables a “multimodal programming” experience (Dietz et al., 2023), in which, users can program through both verbal and visual approaches. Despite the enhanced flexibility of the system, it also causes additional perceptual burden because users need to switch their “brain mode” between “visual thinking” and “text-based thinking”. This finding is aligned with Dual Coding Theory (DCT) (Paivio, 1991), in which Paivio hypothesized that human brains process information using two different channels: verbal and visual. Interestingly, our results seemingly contradict with multiple psychological findings based on DCT that show a combination of verbal and visual information actually helps humans’ memory process (Paivio, 1969; Paivio et al., 2006). For example, people feel it easier to remember a new word if they learn the word using a vocabulary card with a figure with texts. Based on these existing theories on humans’ mind processes, we make the following explanation on the finding in our project: the cause of mental workload does not imply that people dislike a multimodal workspace, but that users require a transparent interface that connects humans’ mental model to the AI reasoning process, both verbally and visually. When users frame the instruction in the InstructPipe dialogue, there is a lack of feedback system from the AI assistant that visualizes how AI interprets the instructions in real time. As a result, humans need to either first estimate the AI prediction or make no expectations when they perform instruction. Even for experts, it is hard to accurately predict AI generation, and thus, the generated pipeline is usually unexpected. When users frame the instructions, the current interface design is purely based on texts, and the visual channel in the brain is off. When the generation is complete, the user suddenly needs to switch on a sleeping (visual) channel, which causes a mental workload. One possible approach to addressing this issue is to expose users to the multimodal information when they perform instructions. For example, future work can investigate methods to visualize a generated pipeline on-the-fly when users are framing and typing their instructions. In this manner, users’ mental models are continuously synchronized with the AI assistant before the final generation, which should eliminate users’ cognitive burdens when the final generation result appears.
  • Debugging. When a rendered pipeline does not match with users’ expectations, users then are required to debug the generated results (see P13’s comment section 6.5.5). In essence, debugging is one professional programming skill, which explains why it mentally overwhelms beginner-level users. InstructPipe made the first step to facilitate users to 1) perceive the results and 2) take actions on the results, i.e., two important elements in a debugging process, on the node-graph editor. Results in our study reveal that this first step is insufficient to support users so that users still feel like they are debugging the codes as professional programmers. Future work should investigate how to further support users’ debugging process with visual programming interfaces. For example, when execution fails, future work can consider visualizing messages that provide cues on the possible issues and provide actionable guidance on the interface for users to fix the issues easily.
On the other hand, it is important to note that we distill the attributes mainly from a subset of participants who explicitly mentioned that they realized an increase in mental workload when using InstructPipe. Those participants typically spent more time writing prompts and working on the node-graph editor than the average performance (e.g., P15 spent the most time ideating and writing the instructions). Therefore, we argue that the aforementioned issues do not represent the average user experience, and the issues are distilled from explicit comments on their negative experience. In fact, we observed that most participants were relatively decisive and could quickly write their instructions and modify a generated result in the node-graph editor. This observation aligns with our qualitative results in Figure 11: the median score of “Mental Demand” in the InstructPipe condition is lower than the baseline condition.

### v2

@@ 6.1. Study Design
In the user evaluation, we aimed to investigate how the interface condition (with InstructPipe and without InstructPipe; the independent variable) affects the user experience and behaviors (the dependent variable). We will refer to these two interface conditions as “InstructPipe” and “Visual Blocks” in the following content. Fig. 7 visualizes the complete study flow. In each condition, participants completed the two pipelines with counterbalance (referred to as Task 1 and Task 2 in Fig. 7).
@@
We carefully designed the experiment to create a fair study that could be completed with reasonable effort. In the following content, we elaborate how we made two important decisions related to the study’s rigor:
6.1.1. Two controlled pipelines with full counterbalancing
Our user evaluation focuses on two controlled pipelines with full counterbalancing. While we acknowledge that more pipelines (e.g., four, six, or more) could enhance generalizability, such designs would also inevitably increase the size of the required user groups, even without fully counterbalancing. For example, fully counterbalancing four controlled pipelines require 12×12$\times$12 × more participants. Partially counterbalancing with four pipelines using the Latin Square design still requires us to double the number of participants. Additionally, novice participants are likely to progressively gain experiences within the study, and such learning effects will weaken the design of partial counterbalancing. We believe that two pipelines with fully counterbalancing is a reasonable experiment setup in this work, and future work could consider extending and scaling up these experiments.
6.1.2. Pipeline selection
Given the fixed number of pipelines we can evaluate with users and the potential bias introduced by few-shot prompts (Zhao et al., 2021a), it is important how we select the two pipelines for user study. There are two critical factors that we considered: representativeness and diversity. Representativeness implies that the selected pipelines should represent the average performance of InstructPipe. Diversity suggests that the selected pipelines should provide various experiences to simulate the actual use scenarios in which the performance of InstructPipe may vary. Following this guideline, we selected four candidates, and the final decision was made after a pilot study with one participant to test the level of pipeline difficulty. The two resulting pipelines are composed of eight nodes with seven edges, and six nodes with six edges, respectively. Using the instructions from two authors, the averaged ratio of human interactions in these two pipelines are $27.8\%$ and $5.2\%$, respectively. See Section D.2 for more detail on the pipelines.
@@ 6.2. Procedure
Each study session takes 55 - 65 minutes in total. The study started with 10-15 minute hands-on training for both conditions. The training included 1) all the Visual Blocks interactions needed to complete the subsequent steps of the experiment, and 2) all the nodes that participants will need to use for pipeline creation in the main session. Participants were also encouraged to experiment with building a pipeline independently, and to ask questions.
After the training, participants progressed to a formal study session where they were asked to build and complete pipelines under the given conditions. We verbally described the pipelines to participants as below, and participants could not see our scripts:
@@
A pipeline is considered complete when the aforementioned functions run in the user’s visual programming workspace. For example, we consider the “real-time multi-model pipeline” as complete when the pipeline registers the sunglasses on the user’s face, with real-time tracking and following of the head movement.
During the task, participants were allowed to consult with us for technical help. If participants were unable to make progress, we provided hints. We provided much more hints in the baseline condition, and we made this decision to ensure every novice-level participant can finish their tasks within a reasonable amount of time. Appendix D.3 contains more details and discussion of the assistance we provided in the study. As an optional extension to the study, eight participants were offered an open-ended pipeline creation, where participants prototyped their own ideas using InstructPipe. This optional section was offered based on the progress of the participant in the previous sections, and time constraints so that the study duration was controlled within the time we guaranteed in our recruitment process.
After conducting all pipeline-condition combinations, participants answered open-ended questions in a semi-structured interview. The interview script is available in the appendix D.1. Participants provided their general impression for each condition, listed pros and cons, identified potential future use cases, and critiqued the user interface for future improvements. We transcribed the interviews and conducted the open coding analysis on the qualitative data (Strauss, 1987; Strauss et al., 1990). More specifically, we categorized the quotes based on our observations and then refined the code for presentation.
@@ 6.3. Participants
We recruited 16 participants from our internal participant pool, which is specifically designed for UX research within our institution. Importantly, none of the participants was involved in our project, and authors in charge of the study did not personally know any of the participants. We screened participants on their self-reported programming experience and machine learning skills. All of the 16 selected participants rated their “Programming Experience” and “’Machine Learning Skill’ as “Intermediate” or below (See Table 3 for a full breakdown). We intentionally recruited novice users, as we envision them as intended users of InstructPipe.
@@ 6.4.1. Task Completion Time
Back-end logs were used to collect timestamps for starting and ending events. Then, completion time for each condition were calculated per task for each participant.
@@ 6.5. Results
Figure 8. Raw-TLX results. The statistic significance is annotated with , ∗∗, or ∗∗∗ (representing $p$¡.05, $p$¡.01, and $p$¡.001, respectively).
[…]
@@ 6.5.1. InstructPipe Reduces Users’ Workload
Fig. 8 further visualizes the results of users’ perceived workload in six sub-scales. The Wilcoxon signed ranks test revealed significant differences on five sub-scales, all except “Mental Demand” (see section 7.2 for more explanations and discussion). Furthermore, the test indicates that all participants unanimously considered that InstructPipe provides lower or equal workload on the subscales of “Physical Demand”, “Temporal Demand”, “Performance” and “Effort” ($W=0$). These quantitative results, with both objective and subjective metrics, demonstrate the potential for InstructPipe to dramatically reduce users’ workload during the pipeline creation process.
@@ 6.5.2. On-boarding Support of Visual Programming
P1, P5, and P9 explicitly mentioned that there is a “learning curve” in visual programming systems, which validates our statements and motivation in section 1.
“There is a learning curve to it (using the visual programming system) for sure, because you have to, like, read each node carefully.” [P1]
P1’s comment matches with our observation of participants’ behaviors during the study. In the Visual Blocks condition, we observed that people were more easily stuck in their creative purposes, which required our support. Typical support included 1) guiding participants if they went too far away from the correct pipeline, and 2) reminding them of an important node for the pipeline, although we introduced all the necessary nodes in our training session.
@@ 6.5.3. Integration into the Existing Workflow
InstructPipe is a feature available in Visual Blocks. In the interviews, participants particularly expressed their strong appreciation of this design as an AI assistant that enhances, instead of completely replacing, the existing user workflow:
@@
While most participants, like P6, appreciated the integration of the AI assistant into the standard visual programming workflow, P15 expressed a concern about this approach. In visual programming, users typically rely on visual thinking to construct pipelines, but the new prompt-based method introduces a shift toward text-based reasoning. This blend of cognitive processes could potentially increase users’ mental workload:
“ [the participant is talking about s/he wants to fix an unsuccessful generation by changing the prompt instead of performing visual programming here] … because I just spent so much time figuring out what the prompt should be. That’s kind of like already where my brain was and I knew that something was wrong there (the prompt), but I would have to switch over to the other mode (visual programming) of figuring out what was wrong in the pipeline … [this is very overwhelming]” [P15]
@@ 6.5.4. Use Scenarios: Accessible ML Prototyping and Education
In the open-ended session, we observed that participants could efficiently utilize InstructPipe to prototype a pipeline for various daily life or business purposes. For example, P14 tried InstructPipe with “summarize real estate price increase in San Diego California over 2023”. Compared to using LLM chatbots, InstructPipe helps the user quickly build a more explainable pipeline in which the user can track (or modify) the information resources. P4 prototyped an interactive VQA app by “Describe the product in the camera”. P13 further shared his thoughts on how this rapid and accessible prototyping experience can support future business:
“It (InstructPipe) is going to facilitate prototype building for PMs (Product Managers) … I have lots of ideas, but my challenge is how to translate an idea into the technical world and see a prototype. I think that this app expedites me in that process a lot.” [P13]
Another emerging theme was regarding educating kids on programming:
@@ 6.5.5. Limitations and Future Directions
Across the study sessions, we consistently observed a specific user behavior pattern: participants typically paused their pace when a generated pipeline appeared in the workspace. At these times, some participants used soliloquy, as in saying “Let me see”, while others kept a focused stare on the workspace. These human behaviors suggest that InstructPipe led participants to engage in deeper, contemplative thought.
The observation suggests that participants needed time to perceive the generated pipelines as they appeared in the workspace. Such sense-making processes bring new challenges to the creative process:
@@
Additionally, we observed that several participants spent more time crafting their prompts than others. P15 spent the most time writing the prompt. The following comments provide insights into how the prompting process caused extra mental workload:
@@ 7.1. Human-AI Collaboration in Prototyping Open-ended ML Pipelines
Our technical evaluation (section 5.3) shows that InstructPipe reduces the number of user interactions to 18.9 % (±$\pm 20.3\%$) There are two key implications from the results:
@@
Such takeaways differ from early-stage findings that show LLMs can achieve full automation of ML inference (Surís et al., 2023; Gupta and Kembhavi, 2023). The main reason is that existing work built their ad hoc solutions for target use scenarios, respectively. In contrast, InstructPipe covers a larger range of ML models (section 3.2) and aims for an open-ended use case. Our results show that LLMs (we used GPT-3.5-turbo in the study) still fail to write robust code with prompt engineering techniques. This aligns with the latest research findings that shows that even the latest LLMs still fail to formulate a whole working pipeline (Wen et al., 2024; Qin et al., 2023b).
While LLMs cannot generate a fully executable pipeline, our work shows that AI can successfully render a certain portion of a pipeline for users. Both technical and user evaluations highlight the important values here. We believe such values provide useful takeaways for HCI researchers to explore more human-AI collaboration approaches and designs in visual programming.
7.2. Three Attributes to Mental Workload
@@
Results in section 6.5.1 show that InstructPipe failed to significantly reduce novice users’ mental demand. We summarized its major causes into three aspects.
7.2.1. Instruction
@@
P15’s comment in section 6.5.5 summarizes the first aspect that causes mental burden. Although the “instruction-to-pipeline” process is fast and seems effortless, the process of framing a prompt is one factor that may overwhelm users, especially those who are more accustomed to visual thinking. InstructPipe requires its users to 1) be clear about the problem they want to solve, and 2) be able to clearly articulate the desired pipeline. Such requirements cause a mental burden to the user. We believe that our results can reinforce Zamfirescu-Pereira et al.’s findings on how non-experts may not prompt LLMs well (Zamfirescu-Pereira et al., 2023) in the visual programming domain.
7.2.2. Perception
@@
The integration of LLM support into the visual programming interface enables a “multimodal programming” experience (Dietz et al., 2023), in which, users can program through both verbal and visual approaches. However, this flexibility increases perceptual burden as users switch between visual and text-based thinking (Paivio, 1991). Interestingly, our results seemingly contradict psychological findings based on the Dual Coding Theory (DCT) that show a combination of verbal and visual information actually helps humans’ memory process [^4] For example, people feel it easier to remember a new word if they learn the word using a vocabulary card with a figure that explains the texts. (Paivio, 1969; Paivio et al., 2006). Therefore, we believe that the mental workload stems not from dislike of multimodal workspaces, but from the lack of a transparent interface that aligns users’ mental models with AI reasoning both verbally and visually. That being said, a next-generation copilot should visualize a pipeline (i.e., visual info) while the user is prompting the system (i.e., verbal info), constituting and interfacing multimodal processing in humans’ brain.
7.2.3. Debugging
When a rendered pipeline does not match users’ expectations, users need to debug (see P13’s comment section 6.5.5). Specifically, users need to “invest extra effort to review and understand the generated content” (Zhang et al., 2023a) and then solve the issues caused by the AI assistant. In essence, debugging is a professional programming skill, which understandably can be mentally overwhelming for beginner-level users. While InstructPipe visualizes generated code in the visual programming platform, our results suggests that future systems should provide more guidance for beginners to better proceed with their programming tasks.